Please help me with this question:
If the probability that a public health nurse will find a client at home is .7, what is the probability (assuming independence) that on two home visits made in a day both clients will be home?
Regards.

Found 2 solutions by ikleyn, ewatrrr:
Answer by ikleyn(53937)

.

For independent events, the probability of two events is the product of individual probabilities


    P(both) = P(one)*P(second) = 0.7*0.7 = 0.49.    ANSWER

Solved, answered, explained and completed.



Answer by ewatrrr(24785)



Hi
p(home) = .7
P(finding one home & the 2nd home) = (.7)(.7) = .49
Wish You the Best in your Studies.